OpenStudy (anonymous):

simplfy rational algebraic expression -6/x+sqrt7

OpenStudy (anonymous):

\[x \sqrt{7} - 6\]

OpenStudy (anonymous):

Cookies pr0x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you sure its not -6+x sqrt7/x?

OpenStudy (anonymous):

\[\frac{-6}{x+\sqrt{7}}\] \[=\frac{-6}{x+\sqrt{7}}\times \frac{x-\sqrt{7}}{x-\sqrt{7}}\] \[=\frac{-6(x-\sqrt{7}}{x^2-7})\]

OpenStudy (anonymous):

or 6 (sqrt7-x)/x^2-7?

OpenStudy (anonymous):

\[-6/x +\sqrt{7}\] ^This?

OpenStudy (anonymous):

i mean the other one..

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

i am willing to bet it is the one i wrote

OpenStudy (anonymous):

ok same ans...thanks i just one to be sure..

OpenStudy (anonymous):

i really dont knw whats the right solution..or procedure..ty again

OpenStudy (anonymous):

\[\frac{(n+1)!}{n!}=n\] is a start

OpenStudy (anonymous):

thanks again...

OpenStudy (anonymous):

forget about the numbers for a minute \[n!=n\times (n-1)\times (n-2)\times ...\times 2\] and \[(n+1)!=(n+1)\times n\times (n-1)\times ...\times 2\]

OpenStudy (anonymous):

so \[\frac{(n+1)!}{n!}\] leads to a whole raft of cancelation, leaving only \[n+1\]

OpenStudy (anonymous):

so all you are left with in your problem is \[\frac{3(n+1)}{5}\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so the ans is x-5/25x^-5?

OpenStudy (anonymous):

@edemsam yes that is it

OpenStudy (anonymous):

ok thanks again..

OpenStudy (anonymous):

but in other site the ans.is 150x^4-30x^2+sqrt(5)(5)/5x^2-1...but they didnt show the steps..